OpenStudy (anonymous):

How do you solve a system where both equations are the equivalent of y? y = 2x – 12 y = -x + 3

OpenStudy (phi):

set their "right hand sides" equal to each other, then simplify

OpenStudy (anonymous):

@Phi How would you do that

OpenStudy (phi):

do you see the right hand side of y = 2x – 12 ? it is 2x -12 set that equal to the right hand side of the other equation

OpenStudy (phi):

can you write down the equation you get when you set the "right hand sides" equal to each other? (Don't solve it, just write it down)

OpenStudy (anonymous):

Oh okay

OpenStudy (phi):

can you write down the equation?

OpenStudy (anonymous):

Yeah 2x-12 -x=3

OpenStudy (phi):

you mean -x+3 on the second line, right? put them on the same line, and put an = between them can you do that ?

OpenStudy (anonymous):

Yeah I meant plus 2x-12=-x+3

OpenStudy (phi):

yes, that looks good. one way to think of this is y = 2x – 12 is (in a way) like saying "the weight of y" is the same as the "weight of 2x - 12" the other equation y = -x + 3 says "the weight of y" is the same as the weight of -x +3 Now we use common sense: if -x + 3 weighs the same as y, and 2x-12 weighs the same as y, then it makes sense that -x+3 weighs the same as 2x-12 we write 2x-12=-x+3

OpenStudy (phi):

to solve, we add numbers (or letters) to both sides (to keep things equal) think of a balance (or scales or a teeter-totter), where you have weight on both sides, and things are balanced.

OpenStudy (phi):

for example add +12 to both sides. Just write +12 on both sides 2x-12=-x+3 add +12 to both sides 2x - 12 + 12 = -x + 3 + 12

OpenStudy (phi):

now simplify the equation 2x - 12 + 12 = -x + 3 + 12 on the right side you have 3+12. can you simplify that to one number ?

OpenStudy (anonymous):

15

OpenStudy (phi):

ok, now you have 2x - 12 + 12 = -x + 15 what about on the left side? what is -12 + 12 ?

OpenStudy (anonymous):

those would cancel out so 0

OpenStudy (anonymous):

so you would be left with 2x=-x+15

OpenStudy (phi):

yes. Now we use the same idea, but with the letters add 1x to both sides 2x + 1x = 1x - 1x + 15 (when they write -x that is the same as -1x) on the left side you have 2 x's and add 1 more x's how many x's do you have ?

OpenStudy (anonymous):

you'd have 3 x's

OpenStudy (phi):

and on the right side we have what ?

OpenStudy (anonymous):

They'd cancel out so we'd be left with 15

OpenStudy (phi):

so now we have 3x = 15 last step. we use the idea that anything divided by itself is 1 Here, we divide 3 by 3. And we have to do the same thing to the other side to keep things balanced 3x/3 = 15/3 on the left 3/3 = 1 1x = 15/3 or (because we don't bother to write the 1 when we have 1x) x = 15/3 simplify the right side, and we get the final answer.

OpenStudy (anonymous):

So x equals 5

OpenStudy (phi):

yes

OpenStudy (anonymous):

Alright thanks

OpenStudy (phi):

yw